OpenStudy (anonymous):

Using the given information, write an equation in slope- intercept form. Then graph. (4,-1) (-2,-2)

OpenStudy (anonymous):

y2-y1 y=mx+b x2-x1 y=-1/6x+b -2-(-1)=-1 (-2)=-1/6(-2)+b -2-4=-6 -2=2/6+b slope=-1/6 subtract 2/6 on both sides -2 2/6 reduced to -2 1/3

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

y=-1/6x-2 1/3

OpenStudy (anonymous):

did this help at all?